# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

### v3

@@ 2.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
$\mathscr{L}_{t}^{\mu}\{f\log f\}$ $=\frac{1}{2}f^{-1}\Gamma_{1}^{t,\mu}(f)+(\log f+1)\mathscr{L}_{t}^{\mu}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),f>0.$
Thus, Itô’s formula and the Kolmogorov backward equation
$\frac{\text{\rm{d}}P_{r,t}^{\mu}f}{\text{\rm{d}}r}=-\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d})$
imply that for $f\in C_{0}^{\infty}(\mathbb{R}^{d})$ with $f>0$,
$\text{\rm{d}}\{(P_{r,t}^{\mu}f)\log(P_{r,t}^{\mu}f)\}(X_{s,r}^{\mu,x})$ $=(-(\log(P_{r,t}^{\mu}f)+1)\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f)(X_{s,r}^{\mu,x})\text{\rm{d}}r$
$\quad+\mathscr{L}_{r}^{\mu}\{(P_{r,t}^{\mu}f)\log(P_{r,t}^{\mu}f)\}(X_{s,r}^{\mu,x})\text{\rm{d}}r+\text{\rm{d}}M_{r}$
$=\frac{1}{2}[(P_{r,t}^{\mu}f)^{-1}[\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)]](X_{s,r}^{\mu,x})\text{\rm{d}}r+\text{\rm{d}}M_{r},\ \ r\in[s,t]$
for some martingale $M_{r}$. Integrating from $s$ to $t$ and taking expectation, we deduce from (2.3) and (2.7) that
$\begin{split}&P_{s,t}^{\mu}(f\log f)(x)-P_{s,t}^{\mu}(f)(x)\log P_{s,t}^{\mu}(f)(x)\\ &=\frac{1}{2}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}[\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)]](X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}|\nabla P_{r,t}^{\mu}f|^{2}](X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}[(P_{r,t}^{\mu}f)^{-1}\text{\rm{e}}^{K_{1}(t-r)}(P_{r,t}^{\mu}|\nabla f|)^{2}](X_{s,r}^{\mu,x})\text{\rm{d}}r.\end{split}$
@@
(2.8) $(P_{r,t}^{\mu}|\nabla f|)^{2}=4\big{(}P_{r,t}^{\mu}\big{(}|\nabla f^{\frac{1}{2}}|\sqrt{f}\big{)}\big{)}^{2}\leq 4P_{r,t}^{\mu}|\nabla f^{\frac{1}{2}}|^{2}P_{r,t}^{\mu}f,$
@@
(2) Inserting $\mu_{t}=P_{t}^{\ast}\mu_{0},t\geq 0$ for $\mu_{0}\in\mathscr{P}_{2}(\mathbb{R}^{d})$, we obtain
(2.9) $(P_{t}^{\ast}\mu_{0})(f):=\int_{\mathbb{R}^{d}}f(x)(P_{t}^{\ast}\mu_{0})(\text{\rm{d}}x)=\int_{\mathbb{R}^{d}}P_{t}^{\mu}f(x)\mu_{0}(\text{\rm{d}}x),\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
where $P_{t}^{\mu}:={P}_{0,t}^{\mu}$. Integrating with respect to $\mu_{0}$ in (2.4) with $s=0$ and applying (2.9), we derive for $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$,
(2.10) $(P_{t}^{\ast}\mu_{0})(f\log f)-\mu_{0}(P_{t}^{\mu}(f)\log P_{t}^{\mu}(f))\leq 2\delta_{1}\frac{\text{\rm{e}}^{K_{1}t}-1}{K_{1}}(P_{t}^{\ast}\mu_{0})\left(|\nabla f^{\frac{1}{2}}|^{2}\right).$
By (2.5), (2.7) and (2.8) for $r=0$ and (2.9), we conclude for $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$,
$\mu_{0}(P_{t}^{\mu}(f)\log P_{t}^{\mu}(f))$ $\leq(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)+2L_{0}\mu_{0}\big{(}|\nabla(\bar{P}_{t}^{\mu_{0}}f)^{\frac{1}{2}}|^{2}\big{)}$
$\leq(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)+2L_{0}\text{\rm{e}}^{K_{1}t}(P_{t}^{\ast}\mu_{0})\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)}.$
This, together with (2.10), implies (2.6) and the proof is completed. ∎
Remark 2.2.
When $\sigma=I_{d\times d}$, the assertions in Theorem 2.1 have been verified in [25, Proposition 1.1]. One can also refer to [8, Theorem 4.1] for the log-Sobolev inequality of time-inhomogeneous semigroup and [29, Lemma 5.2] for the reflecting case under the Bakry-Émery curvature condition. Moreover, (2.7) is equivalent to [8, (10)] if we assume $\sigma\sigma^{\ast}\geq\delta_{2}$ for some $\delta_{2}>0$.
@@
When $\sigma$ depends on the spatial variable and $\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}$ is invertible, we may define a Riemannian metric as
@@
Then the functional inequalities for $P_{s,t}^{\mu}$ follow by using the conclusions in [7] provided that (2.3) holds. However, as explained in [40, p.1451], the assumption (2.3) involves the term $\nabla^{2}[(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast})^{-1}]$, which is more or less implicit. This is the reason why we only consider the case $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$.
Remark 2.4.
$\Gamma^{t,\mu}_{2}(f)\geq\rho\Gamma_{1}^{t,\mu}(f)=\rho|\nabla f|^{2}$
for some constant $\rho\in\mathbb{R}$ are equivalent, see e.g. [22, Proposition 3.12].
2.2 Log-Sobolev inequality for the invariant probability measure of time-homogeneous decoupled SDEs
From now on we will focus on the time-homogeneous version of (1.1), i.e.,
(2.12) $\text{\rm{d}}X_{t}=b(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}W_{t},$
and the associated time-homogeneous decoupled SDE, i.e. for any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, consider
Let $(\tilde{P}_{t}^{\mu})^{\ast}\nu$ be the distribution of the solution to (2.13) with initial distribution $\nu\in\mathscr{P}_{2}(\mathbb{R}^{d})$ and associated time-homogeneous semigroup
$\tilde{P}_{t}^{\mu}f(x)=\int_{\mathbb{R}^{d}}f\text{\rm{d}}\{(\tilde{P}_{t}^{\mu})^{\ast}\delta_{x}\},\ \ f\in\mathscr{B}_{b}(\mathbb{R}^{d}),x\in\mathbb{R}^{d},t\geq 0.$
To establish the log-Sobolev inequality for the invariant probability measure of (2.13), we make use of the following partially dissipative condition.
@@ 2.2 Log-Sobolev inequality for the invariant probability measure of time-homogeneous decoupled SDEs
    $b$ and $\sigma$ are continuous in $\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$. There exist constants $K_{1}>0,K_{2}>0,K_{I}>0,r_{0}>0$ and $\delta_{1}\geq\delta_{2}>0$ such that for any $x,y\in\mathbb{R}^{d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,

### v4

@@ 2.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
(2.15) $\mathscr{L}_{r}^{\mu}\big(g\log g\big)$ $=\frac{1}{2}g^{-1}\Gamma_{1}^{r,\mu}(g)+(\log g+1)\mathscr{L}_{r}^{\mu}g,\ \ g\in C_{b}^{\infty}(\mathbb{R}^{d}),g>0.$
By Itô’s formula, (2.14) and (2.15), we derive that for $f\in C_{b}^{\infty}(\mathbb{R}^{d})$, $f>0$,
$((P_{s^{\prime},t}^{\mu}f)\log(P_{s^{\prime},t}^{\mu}f))(X_{s,s^{\prime}}^{\mu,x})-P_{s,t}^{\mu}f(x)\log P_{s,t}^{\mu}f(x)$
$=\int_{s}^{s^{\prime}}\left(-(\log(P_{r,t}^{\mu}f)+1)\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f+\mathscr{L}_{r}^{\mu}((P_{r,t}^{\mu}f)\log(P_{r,t}^{\mu}f))\right)(X_{s,r}^{\mu,x})\text{\rm{d}}r+M_{s^{\prime}}$
$=\int_{s}^{s^{\prime}}\frac{1}{2}\Big((P_{r,t}^{\mu}f)^{-1}\big(\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f)\big)\Big)(X_{s,r}^{\mu,x})\text{\rm{d}}r+M_{s^{\prime}},\ \ s^{\prime}\in[s,t]$
for some martingale $\{M_{s^{\prime}}\}_{s^{\prime}\in[s,t]}$. Combining this for $s^{\prime}=t$ and taking expectation, we deduce from (2.5) and (2.10) that
$\begin{split}&P_{s,t}^{\mu}(f\log f)(x)-P_{s,t}^{\mu}f(x)\log P_{s,t}^{\mu}f(x)\\ &=\frac{1}{2}\int_{s}^{t}\mathbb{E}\big((P_{r,t}^{\mu}f)^{-1}(\Gamma_{1}^{r,\mu}(P_{r,t}^{\mu}f))\big)(X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}\big((P_{r,t}^{\mu}f)^{-1}|\nabla P_{r,t}^{\mu}f|^{2}\big)(X_{s,r}^{\mu,x})\text{\rm{d}}r\\ &\leq\frac{1}{2}\delta_{1}\int_{s}^{t}\mathbb{E}\big((P_{r,t}^{\mu}f)^{-1}\text{\rm{e}}^{K_{1}(t-r)}(P_{r,t}^{\mu}|\nabla f|)^{2}\big)(X_{s,r}^{\mu,x})\text{\rm{d}}r,\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),\ \ f>0.\end{split}$
@@
(2.16) $(P_{r,t}^{\mu}|\nabla f|)^{2}=4\big(P_{r,t}^{\mu}\big(|\nabla f^{\frac{1}{2}}|f^{\frac{1}{2}}\big)\big)^{2}\leq 4P_{r,t}^{\mu}|\nabla f^{\frac{1}{2}}|^{2}P_{r,t}^{\mu}f,$
@@
Simply denote $P_{t}^{\mu}:={P}_{0,t}^{\mu}$. Integrating with respect to $\gamma$ in (2.6) with $s=0$, we derive for $0<f\in C_{b}^{\infty}(\mathbb{R}^{d})$,
(2.17) $\gamma(P_{t}^{\mu}(f\log f))-\gamma(P_{t}^{\mu}f\log P_{t}^{\mu}f)\leq 2\delta_{1}\frac{\text{\rm{e}}^{K_{1}t}-1}{K_{1}}\gamma\left(P_{t}^{\mu}\left(|\nabla f^{\frac{1}{2}}|^{2}\right)\right).$
By (2.7), (2.10), $P_{r,t}^{\mu}C_{b}^{\infty}(\mathbb{R}^{d})\subset C_{b}^{\infty}(\mathbb{R}^{d})$ and (2.16) for $r=0$, we conclude for $0<f\in C_{b}^{\infty}(\mathbb{R}^{d})$,
$\gamma(P_{t}^{\mu}f\log P_{t}^{\mu}f)$ $\leq\gamma(P_{t}^{\mu}f)\log\gamma(P_{t}^{\mu}f)+2L_{0}\gamma\big(|\nabla(P_{t}^{\mu}f)^{\frac{1}{2}}|^{2}\big)$
$\leq\gamma(P_{t}^{\mu}f)\log\gamma(P_{t}^{\mu}f)+2L_{0}\text{\rm{e}}^{K_{1}t}\gamma\left(P_{t}^{\mu}\left(|\nabla f^{\frac{1}{2}}|^{2}\right)\right).$
This, together with (2.17), implies
(2.18) $\begin{split}&\gamma(P_{t}^{\mu}(f\log f))-\gamma(P_{t}^{\mu}f)\log\gamma(P_{t}^{\mu}f)\\ &\leq 2\Big(L_{0}\text{\rm{e}}^{K_{1}t}+\frac{\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)\Big)\gamma\left(P_{t}^{\mu}\left(|\nabla f^{\frac{1}{2}}|^{2}\right)\right),\ \ 0<f\in C_{b}^{\infty}(\mathbb{R}^{d}),t\geq 0.\end{split}$
(ii) Assume (A), (2.5) and that there exists a locally bounded function $L:[0,\infty)\rightarrow[0,\infty)$ such that for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$,
$|b_{t}(x,\mu_{t})-b_{t}(\bar{x},\mu_{t})|\leq L_{t}|x-\bar{x}|,\ \ t\geq 0,x,\bar{x}\in\mathbb{R}^{d}.$
Let $\hat{b}_{t}^{\mu}(x)=b_{t}(x,\mu_{t})$ and $\hat{\sigma}_{t}^{\mu}=\sigma_{t}(\mu_{t})$ and for any $m\geq 1$, let $\hat{b}^{m,\mu}$ be defined in the same way as $\hat{b}^{m}$ in (4.20) by replacing $\hat{b}$ with $\hat{b}^{\mu}$. Let $P_{s,t}^{m,\mu}$ be the associated semigroup to
$\text{\rm{d}}\hat{X}_{s,t}^{m,\mu}=\hat{b}^{m,\mu}_{t}(\hat{X}_{s,t}^{m,\mu})\text{\rm{d}}t+\hat{\sigma}_{t}^{\mu}\text{\rm{d}}W_{t},\ \ t\geq s\geq 0.$
For $\varepsilon>0$ and $f\in C_{b}^{\infty}(\mathbb{R}^{d})$ with $f>0$, let $f_{\varepsilon}=f+\varepsilon$. By (4.24), (2.9) and (4.2.2) for $\hat{b}^{m,\mu}$ replacing $\hat{b}^{m}$ and step (i), (2.6) and (2.18) hold for $(P_{s,t}^{m,\mu},f_{\varepsilon})$ replacing $(P_{s,t}^{\mu},f)$. Letting $m\rightarrow\infty$, we get (2.6) and (2.18) for $f_{\varepsilon}$ replacing $f$ by the fact $\inf_{x>0}(x\log x)=-\text{\rm{e}}^{-1}$, Fatou’s lemma, (4.26) for $\hat{X}^{m,\mu}$ replacing $\hat{X}^{m}$, the dominated convergence theorem and Lemma 4.3 below for $(\hat{b}^{\mu},\hat{b}^{m,\mu})$ replacing $(\hat{b},\hat{b}^{m})$. Letting $\varepsilon\rightarrow 0$, (2.6) and (2.18) hold by Fatou’s lemma and the monotone convergence theorem.
(iii) Finally, we assume (A) and (2.5). Let $\hat{b}_{t}^{\mu}(x)=b_{t}(x,\mu_{t})$ and $\hat{\sigma}_{t}^{\mu}=\sigma_{t}(\mu_{t})$ and for any $n\geq 1$, let $\hat{b}^{(n),\mu}$ be defined in the same way as $\hat{b}^{(n)}$ in (4.13) by replacing $\hat{b}$ with $\hat{b}^{\mu}$. Let $P_{s,t}^{(n),\mu}$ be the associated semigroup to
$\text{\rm{d}}\hat{X}_{s,t}^{(n),\mu}=\hat{b}^{(n),\mu}_{t}(\hat{X}_{s,t}^{(n),\mu})\text{\rm{d}}t+\hat{\sigma}_{t}^{\mu}\text{\rm{d}}W_{t},\ \ t\geq s\geq 0.$
By (4.14) and (4.16) for $\hat{b}^{(n),\mu}$ replacing $\hat{b}^{(n)}$ and step (ii), (2.6) and (2.18) hold for $(P_{s,t}^{(n),\mu},f_{\varepsilon})$ replacing $(P_{s,t}^{\mu},f)$. Repeating the same argument in step (ii), we obtain (2.6) and (2.18) by Lemma 4.2 in appendix for $(\hat{b}^{\mu},\hat{b}^{(n),\mu})$ replacing $(\hat{b},\hat{b}^{(n)})$. Taking $\mu_{t}=P_{t}^{\ast}\gamma,t\geq 0$ in (2.18) and noting that
$(P_{t}^{\ast}\gamma)(f)=\gamma(P_{t}^{\mu}(f)),\ \ f\geq 0$
and $\inf_{x>0}(x\log x)=-\text{\rm{e}}^{-1}$, we derive (2.8). The proof is completed. ∎
@@
When $\sigma$ depends on the spatial variable, (2.10) does not hold. Instead, if in addition $\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}$ is invertible, one may define a Riemannian metric as
@@
Then the functional inequalities for $P_{s,t}^{\mu}$ follow by using the conclusions in [9] provided that (2.5) holds. However, as explained in [51, p.1451], the assumption (2.5) involves the term $\nabla^{2}\big((\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast})^{-1}\big)$, which is more or less implicit. This is the reason why we only consider the case $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$.
2.2 Time-homogeneous decoupled SDEs
In the following, we consider the time-homogeneous version of (1.1) given by
(2.20) $\text{\rm{d}}X_{t}=b(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}W_{t}.$
We make use of the following partially dissipative condition.
@@ 2.2 Time-homogeneous decoupled SDEs
    $b\colon\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}$ is bounded on bounded sets. For any $\mu\in\mathscr{P}_{2}(\mathbb{R}^{d})$, $b(\cdot,\mu)$ and $\sigma(\cdot,\mu)$ are continuous in $\mathbb{R}^{d}$. There exist constants $K_{1}>0,K_{2}>0,K_{I}>0,r_{0}>0$ and $\delta_{1}\geq\delta_{2}>0$ such that for any $x,y\in\mathbb{R}^{d},\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,